OpenStudy (anonymous):

A candle manufacturer sells cylindrical candles in sets of three. Each candle in the set is a different size. The smallest candle has a radius of 0.5 inches and a height of 3 inches. The other two candles are scaled versions of the smallest, with scale factors of 2 and 3. How much wax is needed to create one set of candles?

OpenStudy (anonymous):

A. 27 B. 36 C. 53 D. 86 E. 98

OpenStudy (anonymous):

@amistre64 @Elise_a18

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

V = pi * r^2 * h V = 3.14 * 0.5^2 * 3 Simplify

OpenStudy (anonymous):

2.355

OpenStudy (elise_a18):

Srry I forget how to do this

OpenStudy (anonymous):

@Elise_a18 its fine

OpenStudy (anonymous):

Yes, now multiply the radius and height by 2 and 3 to find the measurements for the other cylinders. 2 * 0.5 = ? 2 * 3 = ? 3 * 0.5 = ? 3 * 3 = ?

OpenStudy (anonymous):

2 and 3 are the scale factors of the other 2 cylinders.

OpenStudy (anonymous):

1 6 1.5 9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yep, now find the volumes of those. V = pi * r^2 * h V = 3.14 * 1^2 * 6 V = pi * r^2 * h V = 3.14 * 1.5^2 * 9

OpenStudy (anonymous):

@Smileyside

OpenStudy (anonymous):

18.84 63.585

OpenStudy (anonymous):

Yes, now add them all. 2.335 + 18.84 + 63.585

OpenStudy (anonymous):

is that what i was suppose to do?

OpenStudy (anonymous):

2.355*

OpenStudy (anonymous):

=84.76

OpenStudy (anonymous):

Yes, just add them..and whichever option is closest to it is your answer.

OpenStudy (anonymous):

so 86?

OpenStudy (anonymous):

Yep